# Efficient Parameter Optimisation for Quantum Kernel Alignment: A Sub-sampling Approach in Variational Training

### v1

@@ I Introduction
Kernel methods are a fundamental concept in Machine Learning (ML), with a rich history that spans several decades, revolutionising the field by enabling the effective handling of complex, nonlinear relationships in data Hofmann et al. [2008]. The introduction of the “kernel trick" in the 1960s laid the foundation Aiserman et al. [1964], leading to the development of Support Vector Machines (SVMs) in the 1990s Boser et al. [1992], which popularized the concept. Over time, a diverse range of kernel functions have emerged, broadening their utility across many domains. Another milestone was the introduction of the concept of kernel alignment Cristianini et al. [2001]. This method adjusts a kernel function in relation to specific data, thereby limiting the generalisation error Wang et al. [2015], Cortes et al. [2012]. While facing challenges related to scalability Dai et al. [2014], kernel methods continue to be relevant, coexisting with deep learning in solving complex problems.
@@
The field of ML stands to benefit from quantum algorithms, offering the potential for improved scalability and accuracy in solving complex computational tasks Liu et al. [2021], Schuld and Petruccione [2021], Biamonte et al. [2017], Ristè et al. [2017]. A particularly promising approach in this field is the utilization of quantum fidelity kernels, a quantum-enhanced version of kernel methods Schuld and Killoran [2019], Mengoni and Di Pierro [2019], Havlíček et al. [2019], Liu et al. [2021], Huang et al. [2021]. These kernels quantify the overlap between quantum states and leverage the high dimensionality of quantum systems to learn and capture complex patterns in data. A few relevant practical applications of quantum machine learning with kernel methods can be seen in Refs. Mensa et al. [2023], Wu et al. [2021, 2023], Gentinetta et al. [2022, 2023], Miyabe et al. [2023].
@@
Despite the promise of quantum kernel methods, there is scepticism about their practical application due to challenges associated with scalability. These challenges stem from the convergence of quantum kernel values which depends on factors such as data embedding expressibility, the number of qubits, global measurements, entanglement, and noise, which often demand increased measurements for effective training Thanasilp et al. [2022], Cerezo et al. [2022]. In response to these challenges, several techniques have been introduced Huang et al. [2021], Kübler et al. [2021], Shirai et al. [2021], which, along with bandwidth adjustments Canatar et al. [2022], Shaydulin and Wild [2022], aim to enhance generalization in quantum kernels, particularly at higher qubit counts.
@@
Quantum kernel alignment (QKA) adapts classical kernel alignment to utilize quantum kernels, enhancing the model’s precision by aligning more effectively with core data patterns, as outlined in studies by Hubregtsen et al. [2022], Glick et al. [2021], Gentinetta et al. [2023]. Despite this advancement, the scaling in terms of the number of circuits required to align quantum kernels with QKA presents a considerable challenge. In the noiseless case the number of circuits at each training step scales quadratically with dataset size and in the presence of noise it has been shown to scale quartically Gentinetta et al. [2022], which can quickly become prohibitive for larger datasets Miyabe et al. [2023], Thanasilp et al. [2022]. Addressing this complexity with respect to dataset size is crucial for enabling the scalability of QKA methods. One way to address the computational challenges of QKA is to adapt the PEGASOS algorithm Gentinetta et al. [2023]. In reference Gentinetta et al. [2023], they demonstrate that the complexity of this method for an $\epsilon$-accurate classifier with a dataset size of $m$ is $O(min\{m^{2}/\epsilon^{6},1/\epsilon^{1}0\})$ as compared to $O(m^{4.67}/\epsilon^{2})$ for standard quantum kernel alignment in the presence of finite sampling noise. Indeed the PEGASOS-based algorithm scales better than the standard approach with respect to $m$ which makes it attractive for larger datasets but, it comes at the cost of worse scaling with respect to $\epsilon$.
@@
Figure 1: A) Quantum kernel alignment uses n qubits to encode the dataset into a feature map, estimate the kernel, and optimize parameters $\bm{\theta}$ until kernel separation improves. B) In the sub-sampling method, subsets with $k$ elements of the dataset are similarly processed using n qubits. This cycle repeats with different subsets, optimizing $\bm{\theta}$ until the entire dataset is sampled, yielding an optimized kernel. This will end with the estimation of the full dataset, optimised kernel.
II Methodology
@@
Classical kernel methods compute elements of a kernel matrix $K_{x,x^{\prime}}$ using a kernel function $K(\bm{x},\bm{x}^{\prime})$ that measures the distance between two data points $\bm{x}$, $\bm{x}^{\prime}\in X$. For certain choices of distance measure it is possible to use the kernel trick to express the kernel function as an inner product in some vector space $V$. If the kernel function can be written as a feature map $f:X\rightarrow V$, then $V$ is known as the feature space and the kernel function can be written as an inner product on $V$, $K(\bm{x},\bm{x}^{\prime})=\langle f(\bm{x}),f(\bm{x}^{\prime})\rangle_{V}$. We can see from this definition that the kernel function is a measure of similarity between data points that takes on values $[0,1]$ e.g. if $x=x^{\prime}$ then the inner product is one, whereas if the maps $f(x)$, $f(x^{\prime})$ result in orthogonal vectors then the inner product is zero.
@@
Quantum kernel methods can be defined analogously to classical kernel methods. A quantum feature map, $\phi(\bm{x})$, is used to transform classical data $\bm{x}$ into a corresponding quantum state $|\phi(\bm{x})\rangle$. The fidelity kernel between two classical data points, $\bm{x}$ and $\bm{x}^{\prime}$, is computed using the overlap of the related quantum states $K(\bm{x},\bm{x}^{\prime})=|\langle\psi(\bm{x})|\psi(\bm{x}^{\prime})\rangle|^{2}$, where $|\psi(\bm{x})\rangle=U_{\phi}(\bm{x})|0\rangle^{\otimes n}$ and $U_{\phi}(\bm{x})$ is the unitary that implements the quantum feature map. The quality of such kernels can be further improved by quantum kernel alignment techniques that parameterise the kernel and use a variational approach to optimize for the kernel for a specific dataset as suggested in Ref. Hubregtsen et al. [2022], Glick et al. [2021] (see Fig. 1). QKA adds a variational layer $U(\bm{\theta})$ to the kernel, where $\bm{\theta}$ are trainable parameters. The elements of the fidelity kernel matrix are computed in the same way, but now depend on $\bm{\theta}$ as shown in Equation 1,
$K(\bm{x},\bm{x}^{\prime},\bm{\theta})=|\langle\psi(\bm{\theta},\bm{x})|\psi(\bm{\theta},\bm{x}^{\prime})\rangle|^{2},$ (1)
where $|\psi(\bm{\theta},\bm{x})\rangle=U_{\phi}(\bm{x})U(\bm{\theta})|0\rangle^{\otimes n}$.
@@
Similar to classical kernel alignment Cristianini et al. [2001], the training of the variational quantum kernel can be seen as finding a kernel that minimizes the upper bound of the SVM generalization error. This is equivalent to minimising a loss function $L$,
$L(\bm{\theta},a)=\sum_{i}a_{i}-0.5\sum_{i,j}a_{i}a_{j}y_{i}y_{j}K(\bm{\theta},x_{i},x_{j}),$ (2)
@@
where $a_{i}$ and $a_{j}$ are the optimal Lagrange multipliers, and $y$ are the labels of the binary classification problem. The hinge loss, as a function of the variational kernel $K(\bm{\theta})$, is not generally convex, hence the loss landscape may include local minima, complicating the training process.
@@
The full kernel, $K_{f}(\bm{\theta})$ is defined on the full dataset $D$ whereas the sub-kernel $K_{s}(\bm{\theta})$ is defined using a subset of the data $D_{s}\subset D$ that is randomly sampled from $D$. The sub-kernel is used during the variational optimisation phase in order to find optimal parameters $\bm{\theta}_{opt}$. Once obtained, the parameters $\bm{\theta}_{opt}$ are used to construct the full fidelity kernel $K_{f}(\bm{\theta}_{opt})$, which is then used for SVC classification. The benefit of this approach is that fewer circuits are required to build a sub-kernel than the full kernel, therefore training using a sub-kernel can speed-up the training process by a (potentially large) constant factor. The method, schematically represented in Fig. 1 B, can be summarised as follows:
  • Variational parameter initialisation: the variational parameters $\bm{\theta}$ of the quantum circuit are initialised as $\bm{\theta}_{init}$, giving an initial state $|\psi(\bm{\theta}_{init})\rangle$.
@@
    • Loss function optimisation: perform an optimisation step on $L$ to update the parameters $\bm{\theta}$.
  • Full kernel estimation: utilise the optimised parameters $\bm{\theta}_{opt}$ to compute the full quantum kernel $K_{f}(\bm{\theta}_{opt})$.
@@
In the case of training with a full kernel, every iteration involves processing all $m$ data points. The full kernel matrix contains $m^{2}$ entries which correspond to $m^{2}$ evaluations of the kernel function and therefore, assuming no noise, a query complexity of $O(m^{2})$ per training iteration. In the presence of finite sampling noise, it has been shown that sampling overhead results in a considerably worse complexity of $O(m^{4.67}/\epsilon^{2})$ per training iteration for a kernel that is $\epsilon$-close to the infinite shot kernel Gentinetta et al. [2022]. For the remainder of this work, we consider the noise-free case but note that the speed-ups presented may in fact improve in the presence of noise.
Denoting the number of iterations required for convergence as $T$, the overall complexity for the full kernel training is $O(m^{2}T)$. On the other hand, with sub-kernel training, we only process a subset of the data points of size $k$ at each iteration. Given $T^{\prime}$ training iterations and $s$ sub-kernels per iteration, the overall complexity of the sub-sampling approach is $O(sk^{2}T^{\prime})$. As discussed previously, it is likely that $T^{\prime}>T$ meaning it is not guaranteed that the complexity is reduced. It is evident that the upper bound on $T^{\prime}$ is given by $T^{\prime}<\frac{m^{2}}{sk^{2}}T$. Note that, while increasing $s$ decreases the upper bound on $T^{\prime}$, it also decreases the loss error $\epsilon_{s}$ in each iteration which in turn is likely to decrease the actual value of $T^{\prime}$. Assuming that $k<<m$ and that $s<m^{2}/k^{2}$, then $T^{\prime}$ is allowed to be a potentially large multiple of $T$. This opens up the possibility for considerable reductions in run time when training using the sub-sampling approach. While there are no a priori guarantees that $T^{\prime}$ will satisfy the bound given above, our results shown in the next section demonstrate significant speed-ups with little to no loss of classification accuracy relative to the full kernel method.
III Results from Numerical Simulations and Discussion
@@
Our experiments involved two variational ansatze: Real Amplitudes (RA) and Hardware Efficient (HE) Farhi et al. [2014], often referred to as Efficient SU2 ansatze. We used three different optimisers including Gradient Descent (GD) Bottou [2010], SPSA (Simultaneous Perturbation Stochastic Approximation) Spall [1998], ADAM (Adaptive Moment Estimation) Kingma and Ba [2014]. We employed multiple commonly used ansatze and optimisers to test whether or not these choices impact the performance of the sub-sample method. We used different feature maps to encode the data according to the dataset and specify each in the relevant sections. To ensure fairness and reproducibility, the same initial variational parameters are used for every run. Unless stated otherwise, the results presented were obtained from statevector simulations using Qiskit Aleksandrowicz et al. [2019]. The standard deviation of the outcomes through 10-fold cross-validation across all simulations was calculated. We maintained the same test and train subsets for all SVC methods to facilitate an accurate direct comparison between the different algorithms. All experiments were conducted five times, and the results shown are the averages over the values obtained.
@@
In the results presented here, we use the number of queries as a figure of merit. This has the advantage of being a platform-independent metric. Here, a query refers to a single execution of a fidelity calculation i.e., a single circuit. By comparing the number of queries required to reach the optimal solution for each method, we can determine the relative speed-up one method has over another.
III.1 Synthetic Datasets
III.1.1 Benchmarking for Second-order Pauli-Z evolution circuit
@@
For the empirical evaluation of our method, we use a synthetic dataset that is constructed in line with the principles set out by Havlíček et al. Havlíček et al. [2019], where the dataset is specifically designed to be fully separable through a second-order Pauli-Z evolution. The feature map transforms uniformly distributed vectors into quantum states, and labels are attributed based on the inner products of these quantum states. A separation gap of 0.2 is used to set the labeling threshold, ensuring a balanced yet challenging classification problem. The dataset comprises 96 training data points and 32 test data points, each represented by a 2-qubit quantum state.
@@
In Table 1, we show the results obtained from statevector simulation using the SPSA optimiser with different ansatze, sub-samples sizes and the corresponding number of queries and relative speed-up compared to the full kernel evaluation. In this case, we set the maximum number of training iterations to 200 and if the optimiser reaches this point then we pick the point with the lowest loss. If the point with the lowest loss occurs before the final point, we consider this to be the stopping point of the run. All 3 optimisers tested performed comparably, for the full list of results, see the Tables in Appendix B.
Table 1: Results for SPSA optimizer, $k$ represents the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze.
Ansatz k s ROC AUC F1 Queries Speed-up
HE 96 1 0.98 0.931 1.37 ×$\times 10^{6}$ 1.00
HE 32 1 0.99 0.931 3.78 ×$\times 10^{5}$ 3.64
HE 32 4 0.985 0.931 1.38 ×$\times 10^{6}$ 1.01
HE 16 1 0.985 0.856 1.06 ×$\times 10^{5}$ 13.03
HE 16 4 0.99 0.931 4.52 ×$\times 10^{5}$ 3.05
HE 16 8 0.985 0.931 8.95 ×$\times 10^{5}$ 1.54
RA 96 1 0.980 0.931 8.52 ×$\times 10^{5}$ 1.00
RA 32 1 0.985 0.931 3.38 ×$\times 10^{4}$ 25.25
RA 32 8 0.990 0.931 3.06 ×$\times 10^{6}$ 0.28
RA 16 1 0.985 0.931 8.70 ×$\times 10^{4}$ 9.79
RA 16 4 0.985 0.931 2.61 ×$\times 10^{5}$ 3.26
RA 8 8 0.975 0.8986 2.27 ×$\times 10^{5}$ 3.75
The results demonstrate that the sub-sampling method can achieve similar performance using fewer circuits. The number of queries shown in Table 1 is the total number of queries, i.e. the number of queries per iteration multiplied by the number of iterations. We therefore see that, despite the sub-sampling method requiring more iterations to converge the loss function compared to the full kernel method (see Appendix C Figs 5 - 8), the total number of queries is still reduced considerably in almost all cases, resulting in up to x25 speed-up. The only case for which there is a slowdown is for the RA ansatz when $k=32$ and $s=8$. This is perhaps unsurprising given that $m=96$ and so with these parameters the ratio $m^{2}/sk^{2}$ is relatively close to 1. Furthermore, the accuracy of the classification is not deteriorated by the sub-sampling method, showing comparable or even improved accuracy compared to the calculation of the full kernel ($k=96$ and $s=1$ for both HE and RA ansatze).
We also performed a simple hardware experiment for this test case with the SPSA optimiser that can be seen in Appendix B Table 3. The experiments were performed on IBM Quantum hardware, Nairobi, using 2 qubits with 100 shots and no error mitigation. While we cannot draw general conclusions from these experiments, it is interesting to note that the speed-ups are realised even when measuring run time rather than queries. Furthermore, while the performance in terms of ROC AUC and F1 is lower in all cases (full and sub-kernel), we observe several cases where the sub-sampling method improves classification accuracy relative to the full kernel. This is some initial evidence that the method works on real hardware in the presence of noise.
III.1.2 Learning Problem Labeling Cosets with Error (LCE)
The second synthetic dataset we tested was the learning problem Labeling Cosets with Error (LCE) dataset, which was crafted to evaluate the accuracy of QML algorithms. The dataset was generated with a 27-qubit superconducting quantum processor Glick et al. [2021] which is split by 96 training points and 32 test points, each a 7-qubit quantum state defined by Euler angles. These angles specify elements of two separate cosets, forming the basis for binary classification. We apply the same feature map as used in Glick et al. [2021].
@@ III.1.2 Learning Problem Labeling Cosets with Error (LCE)
Table 2: Results for SPSA Optimizer on the Coset dataset, $k$ is the sub-kernel sample size, and $s$ is the number of samples. The F1 and ROC scores are computed on the full, optimised kernel. Results are ordered by descending $k$ and grouped by ansatze.
Ansatz k s ROC AUC F1 Queries Speed-up
HE 96 1 1.00 1.00 1.40 x $10^{6}$ 1.00
HE 32 1 1.00 1.00 3.83 x $10^{5}$ 3.65
HE 32 4 1.00 1.00 1.54 x $10^{6}$ 0.91
HE 16 1 1.00 1.00 1.13 x $10^{5}$ 12.34
HE 16 4 1.00 1.00 4.55 x $10^{5}$ 3.07
HE 16 8 1.00 0.97 9.01 x $10^{5}$ 1.55
RA 96 1 1.00 1.00 1.43 x $10^{6}$ 1.00
RA 32 1 1.00 1.00 3.86 x $10^{5}$ 3.71
RA 32 4 1.00 1.00 1.51 x $10^{6}$ 0.95
RA 16 1 1.00 1.00 1.06 x $10^{5}$ 13.45
RA 16 4 1.00 1.00 4.46 x $10^{5}$ 3.20
RA 16 8 1.00 0.96 9.17 x $10^{5}$ 1.56

### v2

@@ 1 Introduction
Kernel methods are a fundamental concept in Machine Learning (ML), with a rich history that spans several decades, revolutionising the field by enabling the effective handling of complex, nonlinear relationships in data [16]. The introduction of the “kernel trick” in the 1960s laid the foundation [2], leading to the development of Support Vector Machines (SVMs) in the 1990s [4], which popularized the concept. Over time, a diverse range of kernel functions have emerged, broadening their utility across many domains. Another milestone was the introduction of the concept of kernel alignment [9]. This method adjusts a kernel function in relation to specific data, thereby limiting the generalisation error [35, 8]. While facing challenges related to scalability [10], kernel methods continue to be relevant, coexisting with deep learning in solving complex problems.
@@
The field of ML stands to benefit from quantum algorithms, offering the potential for improved scalability and accuracy in solving complex computational tasks [22, 30, 3, 28]. A particularly promising approach in this field is the utilization of quantum fidelity kernels, a quantum-enhanced version of kernel methods [29, 23, 15, 22, 17]. These kernels quantify the overlap between quantum states and leverage the high dimensionality of quantum systems to learn and capture complex patterns in data. A few relevant practical applications of quantum machine learning with kernel methods can be seen in Refs. [24, 36, 37, 13, 12, 25].
@@
Despite the promise of quantum kernel methods, there is scepticism about their practical application due to challenges associated with scalability. These challenges stem from the convergence of quantum kernel values which depends on factors such as data embedding expressibility, the number of qubits, global measurements, entanglement, and noise, which often demand increased measurements for effective training [34, 7]. In response to these challenges, several techniques have been introduced [17, 21, 32], which, along with bandwidth adjustments [6, 31], aim to enhance generalization in quantum kernels, particularly at higher qubit counts.
@@
Quantum kernel alignment (QKA) adapts classical kernel alignment to utilize quantum kernels, enhancing the model’s precision by aligning more effectively with core data patterns, as outlined in studies by [18, 14, 12]. Despite this advancement, the scaling in terms of the number of circuits required to align quantum kernels with QKA presents a considerable challenge. In the noiseless case the number of circuits at each training step scales quadratically with dataset size and in the presence of noise it has been shown to scale quartically [13], which can quickly become prohibitive for larger datasets [25, 34]. Addressing this complexity with respect to dataset size is crucial for enabling the scalability of QKA methods. One way to address the computational challenges of QKA is to adapt the PEGASOS algorithm [12]. In reference [12], they demonstrate that the complexity of this method for an $\epsilon$-accurate classifier with a dataset size of $m$ is $O(min\{m^{2}/\epsilon^{6},1/\epsilon^{10}\})$ as compared to $O(m^{4.67}/\epsilon^{2})$ for standard quantum kernel alignment in the presence of finite sampling noise. Indeed the PEGASOS-based algorithm scales better than the standard approach with respect to $m$ which makes it attractive for larger datasets but, it comes at the cost of worse scaling with respect to $\epsilon$.
@@
Figure 1: A) Quantum kernel alignment uses n qubits to encode the dataset into a feature map, estimate the kernel, and optimize parameters $\boldsymbol{\theta}$ until kernel separation improves. B) In the sub-sampling method, subsets with $k$ elements of the dataset are similarly processed using n qubits. This cycle repeats with different subsets, optimizing $\boldsymbol{\theta}$ until the entire dataset is sampled, yielding an optimized kernel. This will end with the estimation of the full dataset, optimised kernel.
2 Methodology
@@
Classical kernel methods compute elements of a kernel matrix $K_{x,x^{\prime}}$ using a kernel function $K(\boldsymbol{x},\boldsymbol{x}^{\prime})$ that measures the distance between two data points $\boldsymbol{x}$, $\boldsymbol{x}^{\prime}\in X$. For certain choices of distance measure it is possible to use the kernel trick to express the kernel function as an inner product in some vector space $V$. If the kernel function can be written as a feature map $f:X\rightarrow V$, then $V$ is known as the feature space and the kernel function can be written as an inner product on $V$, $K(\boldsymbol{x},\boldsymbol{x}^{\prime})=\langle f(\boldsymbol{x}),f(\boldsymbol{x}^{\prime})\rangle_{V}$. We can see from this definition that the kernel function is a measure of similarity between data points that takes on values $[0,1]$ e.g. if $x=x^{\prime}$ then the inner product is one, whereas if the maps $f(x)$, $f(x^{\prime})$ result in orthogonal vectors then the inner product is zero.
@@
Quantum kernel methods can be defined analogously to classical kernel methods. A quantum feature map, $\phi(\boldsymbol{x})$, is used to transform classical data $\boldsymbol{x}$ into a corresponding quantum state $|\phi(\boldsymbol{x})\rangle$. The fidelity kernel between two classical data points, $\boldsymbol{x}$ and $\boldsymbol{x}^{\prime}$, is computed using the overlap of the related quantum states $K(\boldsymbol{x},\boldsymbol{x}^{\prime})=|\langle\psi(\boldsymbol{x})|\psi(\boldsymbol{x}^{\prime})\rangle|^{2}$, where $|\psi(\boldsymbol{x})\rangle=U_{\phi}(\boldsymbol{x})|0\rangle^{\otimes n}$ and $U_{\phi}(\boldsymbol{x})$ is the unitary that implements the quantum feature map. The quality of such kernels can be further improved by quantum kernel alignment techniques that parameterise the kernel and use a variational approach to optimize for the kernel for a specific dataset as suggested in Refs. [18, 14] (see Fig. 1). QKA adds a variational layer $U(\boldsymbol{\theta})$ to the kernel, where $\boldsymbol{\theta}$ are trainable parameters. The elements of the fidelity kernel matrix are computed in the same way, but now depend on $\boldsymbol{\theta}$ as shown in Equation 1,
$K(\boldsymbol{x},\boldsymbol{x}^{\prime},\boldsymbol{\theta})=|\langle\psi(\boldsymbol{\theta},\boldsymbol{x})|\psi(\boldsymbol{\theta},\boldsymbol{x}^{\prime})\rangle|^{2},$ (1)
where $|\psi(\boldsymbol{\theta},\boldsymbol{x})\rangle=U_{\phi}(\boldsymbol{x})U(\boldsymbol{\theta})|0\rangle^{\otimes n}$. Note that, while we focus on kernels of the form shown in Equation (1), in principle our method generalises to other choices of kernel.
@@
Similar to classical kernel alignment [9], the training of the variational quantum kernel can be seen as finding a kernel that minimizes the upper bound of the SVM generalization error. This is equivalent to minimising a loss function $L$,
$L(\boldsymbol{\theta},a)=\sum_{i}a_{i}-0.5\sum_{i,j}a_{i}a_{j}y_{i}y_{j}K(\boldsymbol{\theta},x_{i},x_{j}),$ (2)
@@
where $a_{i}$ and $a_{j}$ are the optimal Lagrange multipliers, and $y$ are the labels of the binary classification problem. The hinge loss, as a function of the variational kernel $K(\boldsymbol{\theta})$, is not generally convex, hence the loss landscape may include local minima, complicating the training process.
@@
The full kernel, $K_{f}(\boldsymbol{\theta})$ is defined on the full dataset $D$ whereas the sub-kernel $K_{s}(\boldsymbol{\theta})$ is defined using a subset of the data $D_{s}\subset D$ that is randomly sampled from $D$. The sub-kernel is used during the variational optimisation phase in order to find optimal parameters $\boldsymbol{\theta}_{opt}$. Once obtained, the parameters $\boldsymbol{\theta}_{opt}$ are used to construct the full fidelity kernel $K_{f}(\boldsymbol{\theta}_{opt})$, which is then used by a Support Vector Classifier (SVC). The benefit of this approach is that fewer circuits are required to build a sub-kernel than the full kernel, therefore training using a sub-kernel can speed-up the training process by a factor that is at worst constant and at best removes the dependence of scaling on $m$. The method, schematically represented in Fig. 1 B, can be summarised as follows:
  • Variational parameter initialisation: the variational parameters $\boldsymbol{\theta}$ of the quantum circuit are initialised as $\boldsymbol{\theta}_{init}$, giving an initial state $|\psi(\boldsymbol{\theta}_{init})\rangle$.
@@
    • Loss function optimisation: perform an optimisation step on $L$ to update the parameters $\boldsymbol{\theta}$.
  • Full kernel estimation: utilise the optimised parameters $\boldsymbol{\theta}_{opt}$ to compute the full quantum kernel $K_{f}(\boldsymbol{\theta}_{opt})$.
@@
In the case of training with a full kernel, every iteration involves processing all $m$ data points. The full kernel matrix contains $m^{2}$ entries which correspond to $m^{2}$ evaluations of the kernel function and therefore, assuming no noise, a query complexity of $O(m^{2})$ per training iteration. In the presence of finite sampling noise, it has been shown that sampling overhead results in a considerably worse complexity of $O(m^{4.67}/\epsilon^{2})$ per training iteration for a kernel that is $\epsilon$-close to the infinite shot kernel [13]. For the remainder of this work, we consider the noise-free case but note that the speed-ups presented may in fact improve in the presence of noise.
Denoting the number of iterations required for convergence as $T$, the overall complexity for the full kernel training is $O(m^{2}T)$. On the other hand, with sub-kernel training, we only process a subset of the data points of size $k$ at each iteration. Given $T^{\prime}$ training iterations and $s$ sub-kernels per iteration, the overall complexity of the sub-sampling approach is $O(sk^{2}T^{\prime})$. As discussed previously, it is likely that $T^{\prime}>T$ meaning it is not guaranteed that the complexity is reduced. It is evident that the upper bound on $T^{\prime}$ is given by $T^{\prime}<\frac{m^{2}}{sk^{2}}T$. Note that, while increasing $s$ decreases the upper bound on $T^{\prime}$, it also decreases the loss error $\epsilon_{s}$ in each iteration which in turn is likely to decrease the actual value of $T^{\prime}$. Assuming that $k<<m$ and that $s<m^{2}/k^{2}$, then $T^{\prime}$ is allowed to be a potentially large multiple of $T$. This opens up the possibility for considerable reductions in run time when training using the sub-sampling approach. Note that, if $s$ and $k$ are chosen to be constants independent of $m$ and $T^{\prime}$ does not scale with $m$, then the overall complexity of the training process is independent of $m$. While there are no a priori guarantees that $T^{\prime}$ will satisfy the bound given above, our results shown in the next section demonstrate significant speed-ups with little to no loss of classification accuracy relative to the full kernel method for choices of $s$ and $k$ that are independent of $m$.
3 Results from Numerical Simulations and Discussion
@@
Our experiments involved two variational ansatze: Real Amplitudes (RA) and Hardware Efficient (HE) [11], often referred to as Efficient SU2 ansatze. We used three different optimisers including Gradient Descent (GD) [5], SPSA (Simultaneous Perturbation Stochastic Approximation) [33], ADAM (Adaptive Moment Estimation) [20]. We employed multiple commonly used ansatze and optimisers to test whether or not these choices impact the performance of the sub-sample method. We used different feature maps to encode the data according to the dataset and specify each in the relevant sections. To ensure fairness and reproducibility, the same initial variational parameters are used for every run. Unless stated otherwise, the results presented were obtained from statevector simulations using Qiskit [19]. The standard deviation of the outcomes through 10-fold cross-validation across all simulations was calculated. We maintained the same test and train subsets for all SVC methods to facilitate an accurate direct comparison between the different algorithms. All experiments were conducted five times, and the results shown are the averages over the values obtained.
@@
In the results presented here, we use the number of queries as a figure of merit. This has the advantage of being a platform-independent metric. Here, a query refers to a single execution of a fidelity calculation i.e., a single circuit. By comparing the number of queries required to reach the optimal solution for each method, we can determine the relative speed-up one method has over another. Specifically, we define speed-up to be the ratio of the number of queries required for a full kernel compared to a given sub-kernel of size $k$.
3.1 Synthetic Datasets
3.1.1 Benchmarking for Second-order Pauli-Z evolution circuit
@@
For the empirical evaluation of our method, we use a synthetic dataset that is constructed in line with the principles set out by Havlíček et al. [15], where the dataset is specifically designed to be fully separable through a second-order Pauli-Z evolution. The feature map transforms uniformly distributed vectors into quantum states, and labels are attributed based on the inner products of these quantum states. A separation gap of 0.2 is used to set the labeling threshold, ensuring a balanced yet challenging classification problem. The dataset comprises 96 training data points and 32 test data points, each represented by a 2-qubit quantum state.
@@
In Fig. 2, we show the results obtained from statevector simulation using the SPSA optimiser with different ansatze, sub-samples sizes and the corresponding number of queries and relative speed-up compared to the full kernel evaluation. In this case, we set the maximum number of training iterations to 200 and if the optimiser reaches this point then we pick the point with the lowest loss. If the point with the lowest loss occurs before the final point, we consider this to be the stopping point of the run. In practice, a different stopping criterion such as a loss threshold could be chosen to avoid running more iterations than necessary. All 3 optimisers tested performed comparably, for the full list of results, see the Tables in Appendix B.
[…]
@@
Figure 2: Total number of queries and speed-up relative to full kernel for the top 12 results using the SPSA optimizer, grouped by Hardware Efficient (HE) ansatze and Real Amplitudes (RA) ansatze. Upper panels: the range of ROC AUC values is 0.975-0.990 and the F1 scores range from 0.856-0.931 on the ZZ dataset (see Table 3 for full results). Lower panels: all ROC AUC values are 1.0 and the F1 values range from 0.969-1.0 on the Coset dataset (see Table 7 for full results). In both panels, $k$ is the sub-kernel size, and $s$ is the number of samples. Results are ordered by descending $k$.
For the HE Ansatz results shown in the left panel of Fig. 2, the full kernel achieves an ROC AUC of 0.98 and an F1 score of 0.93. For all choices of $k$ and $s$, the ROC AUC is slightly improved relative to the full kernel, with an average improvement of 0.7%. The F1 score achieved is identical to the full kernel in all cases apart from $k=16$, $s=1$ which sees an 8.2% drop in F1 score. For the RA Ansatz results shown in the right panel of Fig. 2, the full kernel achieves the same ROC AUC and F1 score as for the HE Ansatz. The parameters $k=8$ and $s=8$ are the only choice for which the ROC AUC and F1 decrease relative to the full kernel (by 0.51% and 3.5% respectively). For all other parameter choices shown the F1 score is identical to the full kernel score and the ROC AUC increases by 0.51-1.02%.
The results demonstrate that the sub-sampling method can achieve similar performance using fewer circuits. The number of queries shown in Figure 2 is the total number of queries, i.e. the number of queries per iteration multiplied by the number of iterations. We therefore see that, despite the sub-sampling method requiring more iterations to converge the loss function compared to the full kernel method (see Appendix D Figs 6 - 9), the total number of queries is still reduced considerably in almost all cases, resulting in up to x25 speed-up (reduction in number of queries). The only case for which there is a slowdown is for the RA ansatz when $k=32$ and $s=8$. This is perhaps unsurprising given that $m=96$ and so with these parameters the ratio $m^{2}/sk^{2}$ is relatively close to 1. Furthermore, the accuracy of the classification is not deteriorated by the sub-sampling method, showing comparable or even improved accuracy compared to the calculation of the full kernel ($k=96$ and $s=1$ for both HE and RA ansatze).
We also performed a simple hardware experiment for this test case with the SPSA optimiser that can be seen in Appendix C Table 2. The experiments were performed on IBM Quantum hardware, Nairobi, using 2 qubits with 100 shots and no error mitigation. While we cannot draw general conclusions from these experiments, it is interesting to note that the speed-ups are realised even when measuring run time rather than queries. Furthermore, while the performance in terms of ROC AUC and F1 is lower in all cases (full and sub-kernel), we observe several cases where the sub-sampling method improves classification accuracy relative to the full kernel. This is some initial evidence that the method works on real hardware in the presence of noise.
3.1.2 Learning Problem Labeling Cosets with Error (LCE)
The second synthetic dataset we tested was the learning problem Labeling Cosets with Error (LCE) dataset, which was crafted to evaluate the accuracy of QML algorithms. The dataset was generated with a 27-qubit superconducting quantum processor [14] which is split by 96 training points and 32 test points, each a 7-qubit quantum state defined by Euler angles. These angles specify elements of two separate cosets, forming the basis for binary classification. We apply the same feature map as used in [14].